OpenStudy (anonymous):

Solve the equations and check your solutions. If there is no solution, say no 4x/x+1=4x-3/x-2

OpenStudy (anonymous):

Does it look like this: \(\dfrac{4x}{x+1}= \dfrac{4x-3}{x-2}\)

OpenStudy (anonymous):

If so, start by cross multiplying.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4x/2(x+1)-4x-3/x-2 4x/x+1-4x-3/x-2=0 4x(x0^2-4x-3(x+1)/x+1(x-2)x-2(x+1) 4x^2-8x-4x6+x-5

OpenStudy (anonymous):

FYI: Using the equation editor will make your equations more clear... I don't think you got the cross multiplication correct. Here is an example. \(\dfrac{x}{y} = \dfrac{a}{b}\) \(xb=ya\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

got figure out how to use the equation editor

OpenStudy (anonymous):

Your equation should look like this after you cross multiply: \(4x^2-8x=4x^2-3x+4x-3\)

OpenStudy (anonymous):

This site uses \(LaTeX\), I personally find that much easier to use. To see the code that someone else used, right click \(\rightarrow\) "Show Math As" \(\rightarrow\) "Tex Commands".

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\(4x^2-8x=4x^2-3x+4x-3\) Is written as: `\(4x^2-8x=4x^2-3x+4x-3\)`

OpenStudy (anonymous):

Back to your question, do you see how I did the cross multiplication?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great! Let me know if you need any more help.

OpenStudy (anonymous):

\(4x^2-8x=4x^2-3x+4x-3/)

OpenStudy (anonymous):

You're slash on the right end is facing the wrong direction... \(4x^2-8x=4x^2-3x+4x-3\)

OpenStudy (anonymous):

My grammar/spelling gets bad after 11pm. Sorry.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

From here, combine like terms.

OpenStudy (anonymous):

4x^2-8x=16x^2-3x+3

OpenStudy (anonymous):

Use inverse operations: addition<->subtraction: \(\cancel{4x^2-4x^2}-8x=\cancel{4x^2-4x^2}-3x+4x-3\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

there is no solution